# Non-Convex Stochastic Composite Optimization with Polyak Momentum

### v1

@@ 5 The algorithm and analysis
In this section, we present the stochastic proximal gradient method with Polyak momentum and give its convergence guarantees in the non-convex regime.
@@
6:$\Omega_{k}(\mathbf{x})\coloneqq\langle\mathbf{m}_{k},\mathbf{x}\rangle+\psi(\mathbf{x})+\frac{M}{2}\left\lVert\mathbf{x}-\mathbf{x}_{k}\right\rVert^{2}$
7:s.t. $\Omega_{k}(\mathbf{x}_{k+1})\leq\Omega_{k}(\mathbf{x}_{k})$ and $\left\lVert\nabla\Omega_{k}(\mathbf{x}_{k+1})\right\rVert^{2}\leq\delta_{k}$
@@
The algorithm is given in Algorithm 1. In the algorithm, the stochastic gradient $\mathbf{g}_{k}$ at each iteration $k$ is replaced by the momentum - aggregated gradient mean:
$\mathbf{m}_{k}=(1-\gamma)\mathbf{m}_{k-1}+\gamma\mathbf{g}_{k}\,.$
@@
We relax 5.1 and discuss the inexact proximal steps in Section 7.
@@ 5.1 Convergence analysis
We discuss the convergence analysis here, and missing proofs can be found in Appendix A. The convergence analysis of Algorithm 1 revolves around the following quantities:
@@
$F_{k}$ quantifies the distance between the current objective value and the lower bound. $\Delta_{k}$ bounds the distance between the current gradient estimate and the full gradient. $R_{k}$ bounds the distance between two consecutive iterates. We start by giving a descent lemma on $\Delta_{k}$, which is key to analyzing the variance reduction effect of momentum. Similar statements can be found in [7] and [12].
Lemma 5.2.
Given 3.4, and for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, we have:
$\Delta_{k+1}\leq(1-\gamma)\Delta_{k}+\frac{L^{2}}{\gamma}R_{k}+\gamma^{2}\sigma^{2}\,.$
The $(1-\gamma)$ factor in the above lemma is the key to show that $\Delta_{k}$ decreases in the optimization process. Next, we discuss the per-iteration descent of $F_{k}$.
Recall that in this section, we have 5.1, which is equivalent to the following stationarity assumption:
$\nabla\Omega_{k}(\mathbf{x}_{k+1})=\nabla\psi(\mathbf{x}_{k+1})+\mathbf{m}_{k}+M(\mathbf{x}_{k+1}-\mathbf{x}_{k})=0\,.$ (5)
@@
Lemma 5.3.
Given 3.1, 3.3, and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, we have:
$F_{k+1}\leq F_{k}-\frac{M-L}{4}R_{k}+\frac{\Delta_{k}}{M-L}\,.$
Remark 5.4.
The constants in Lemma 5.3 can be slightly improved if 3.1 is replaced by the convexity of $\psi$.
@@
$\leq f(\mathbf{x}_{k})+\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle+\frac{L}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+\psi(\mathbf{x}_{k+1})$
$=f(\mathbf{x}_{k})+\langle\mathbf{m}_{k},\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle+\psi(\mathbf{x}_{k+1})+\frac{M}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}$
$\quad-\frac{M-L}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+\langle\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k},\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle\,.$
Then we apply 3.1 and notice that $\Omega_{k}(\mathbf{x}_{k})=\psi(\mathbf{x}_{k})$:
$F(\mathbf{x}_{k+1})$ $\leq F(\mathbf{x}_{k})-\frac{M-L}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+\langle\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k},\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle\,.$
@@
$F(\mathbf{x}_{k+1})$ $\leq F(\mathbf{x}_{k})-\frac{M-L}{4}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+\frac{\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}}{M-L}\,.$
@@
Lemma 5.5.
Given 3.3 and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, we have:
$R_{k}\geq\frac{{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]}{3(M^{2}+L^{2})}-\frac{\Delta_{k}}{M^{2}+L^{2}}\,.$
@@
Note that we can split $\nabla F(\mathbf{x}_{k+1})$ in the following way:
@@
$=\mathbf{m}_{k}+\nabla\psi(\mathbf{x}_{k+1})+(\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k})+(\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k}))\,.$
Then we have:
@@
Now apply the stationarity condition (5) on the first term, and use 3.3 on the third term, we get:
$\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}$ $\leq 3(M^{2}+L^{2})\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+3\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}\,.$
We get the desired result by rearranging and taking expectations on both sides. ∎
@@
Lemma 5.6.
Given 3.1, 3.3, 3.4, and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\frac{3L}{(M-L)},a\coloneqq\frac{3}{8L}$, and $M>4L$, we have:
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{27L}{4M^{2}}\sigma^{2}\,.$ (7)
Remark 5.7.
Note that the above lemma sets the momentum parameter $\gamma$ to be of the order $\Theta(\frac{L}{M})$.
We defer the proof to Appendix A. Now, with the above lemma and following the standard arguments, for a total number of $K$ iterations, and for $M=4L+\frac{3^{\nicefrac{{3}}{{2}}}}{2}\sqrt{\frac{KL\sigma^{2}}{\Phi_{0}}}$, we have:
$\frac{1}{K}\sum_{k=0}^{K-1}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]\leq\mathcal{O}\left(\sqrt{\frac{L\Phi_{0}\sigma^{2}}{K}}+\frac{L\Phi_{0}}{K}\right)\,.$
This indicates that the final convergence guarantee depends on $\Phi_{0}$, i.e. the initialization, which we discuss in the next section.
@@ 5.2 Initialization and convergence guarantees
As mentioned in Remark 5.7, $a=\mathcal{O}(\frac{1}{L})$, therefore $\Phi_{0}=\mathcal{O}(F_{0}+\frac{\Delta_{0}}{L})$. There are some subtle differences in how $\Delta_{0}$ can be initialized between the non-composite and composite cases.
@@
When $\psi\equiv 0$, i.e. $F\equiv f$, we set the initial momentum $\mathbf{m}_{-1}\coloneqq-\frac{\gamma\mathbf{g}_{0}}{1-\gamma}$, then $\Delta_{0}={\mathbb{E}}\left[\left\lVert\mathbf{m}_{0}-\nabla f(\mathbf{x}_{0})\right\rVert^{2}\right]=\left\lVert\nabla f(\mathbf{x}_{0})\right\rVert^{2}\leq\mathcal{O}(LF_{0})$, where the last inequality follows from 3.3 and the fact that $F\equiv f$. Therefore, in the non-composite case, we can initialize the parameters such that $\Phi_{0}=\mathcal{O}(F_{0})$.
@@
The composite case is slightly trickier. We set $\mathbf{m}_{-1}\coloneqq\mathbf{g}_{0}$ and get that $\mathbf{m}_{0}=\mathbf{g}_{0}$. Hence $\Delta_{0}={\mathbb{E}}\left[\left\lVert\mathbf{g}_{0}-\nabla\psi(\mathbf{x}_{0})\right\rVert^{2}\right]\leq\sigma^{2}$. Therefore, $\Phi_{0}=\mathcal{O}(F_{0}+\frac{\sigma^{2}}{L})$. When $\sigma^{2}=\mathcal{O}(LF_{0})$, we get the same $\Phi_{0}=\mathcal{O}(F_{0})$ as in the non-composite case.
@@
In the case that $\sigma^{2}$ is much larger than $F_{0}$, if we have access to a constant size (not depending on the target error) mini-batch initially, then we can set $\mathbf{g}_{0}=\frac{1}{b_{0}}\sum_{i=1}^{b_{0}}\nabla f(\mathbf{x}_{0},\xi_{i})$ where $b_{0}\coloneqq\frac{\sigma^{2}}{LF_{0}}$, i.e. $\mathbf{g}_{0}$ is a mini-batch stochastic gradient of batch size $\frac{\sigma^{2}}{LF_{0}}$. Then we have $\Phi_{0}=\mathcal{O}(F_{0}+\frac{\sigma^{2}}{b_{0}L})=\mathcal{O}(F_{0})$, which is the same as in the non-composite case.
Now we state the final convergence guarantees of Algorithm 1:
Theorem 5.8.
Given 3.1, 3.3, 3.4, and 5.1, then for $\{\mathbf{x}_{k}\}_{k>0}$ generated by Algorithm 1, if $\gamma\coloneqq\frac{3L}{(M-L)}$, there exists $M=4L+\frac{3^{\nicefrac{{3}}{{2}}}}{2}\sqrt{\frac{KL\sigma^{2}}{\Phi_{0}}}$ such that after at most
$K=\mathcal{O}\left(\frac{L\Phi_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{L\Phi_{0}}{\varepsilon}\right)$
iterations of Algorithm 1 it holds that ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{\kappa})\right\rVert^{2}\right]\leq\varepsilon$, where $\mathbf{x}_{\kappa}$ is chosen from $\kappa\in\left\{1,\cdots,K\right\}$ uniformly at random.
More specifically, if $\psi\equiv 0$, we can initialize $\mathbf{m}_{-1}$ such that $K=\mathcal{O}(\frac{LF_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{LF_{0}}{\varepsilon})$; Otherwise, we can initialize $\mathbf{m}_{-1}$ such that $K=\mathcal{O}(\frac{LF_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{\sigma^{4}}{\varepsilon^{2}}+\frac{LF_{0}}{\varepsilon}+\frac{\sigma^{2}}{\varepsilon})$; further, when an error-independent initial mini-batch is allowed, we can initialize $\mathbf{m}_{-1}$ with a $\frac{\sigma^{2}}{LF_{0}}$ batch such that $K=\mathcal{O}(\frac{LF_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{LF_{0}}{\varepsilon})$.
We note that the output criteria (selecting a random iterate) is standard in the literature and could also be replaced with an average of all iterates. In both cases, this can tracked efficiently (e.g. [33]).
When an initial mini-batch is not allowed, the convergence rate has a $\mathcal{O}(\frac{\sigma^{4}}{\varepsilon^{2}}+\frac{\sigma^{2}}{\varepsilon})$ term that is not present in the non-composite case. One natural question is whether this is an artifact of our analysis or inherent to the problem and the algorithm. Note that the first step of the Algorithm 1 coincides with the vanilla stochastic proximal gradient method. Consider the lower bound construction in Proposition 4.1: when starting at $\mathbf{x}_{0}=0$, which is a stationary point such that $F_{0}=0$, one such step would incur an $\mathcal{O}(\sigma^{2})$ error. Therefore, in the non-composite case without the initial mini-batch, the convergence rate must have some term that only depends on $\sigma^{2}$. In other words, the extra terms seem unavoidable.

### v4

@@ 5 The Algorithm and Analysis
In this section, we present the stochastic proximal gradient method with the Polyak momentum and establish its convergence guarantees in the non-convex regime.
@@
The method is shown in Algorithm 1. In the algorithm, the stochastic gradient $\mathbf{g}_{k}$ at each iteration $k$ is replaced by the momentum-aggregated gradient mean:
$\mathbf{m}_{k}=(1-\gamma_{k-1})\mathbf{m}_{k-1}+\gamma_{k-1}\mathbf{g}_{k}\,.$
@@
We relax Assumption 5 and discuss the inexact proximal steps in Section 7.
@@ 5.1 Convergence Analysis
We discuss the convergence analysis here, and missing proofs can be found in Appendix B. The convergence analysis of Algorithm 1 revolves around the following quantities:
@@
$F_{k}$ quantifies the distance between the current objective value and the lower bound. $\Delta_{k}$ bounds the distance between the current gradient estimate and the full gradient. $R_{k}$ bounds the distance between two consecutive iterates. We start by giving a descent lemma on $\Delta_{k}$, which is key to analyzing the variance reduction effect of momentum. Similar statements can be found in Cutkosky and Mehta (2020) and Fatkhullin et al. (2023).
Lemma 1.
Under Assumption 4, for any $k\geq 0$:
$\Delta_{k+1}\leq(1-\gamma_{k})\Delta_{k}+\frac{L^{2}}{\gamma_{k}}R_{k}+\gamma_{k}^{2}\sigma^{2}\,.$
The $(1-\gamma_{k})$ factor in the above lemma is the key to show that $\Delta_{k}$ decreases in the optimization process. Next, we discuss the per-iteration descent of $F_{k}$.
Recall that in this section, we have Assumption 5, which is equivalent to the following stationarity assumption:
$\nabla\psi(\mathbf{x}_{k+1})+\mathbf{m}_{k}+M_{k}(\mathbf{x}_{k+1}-\mathbf{x}_{k})=0.$ (5)
@@
Lemma 2.
Under Assumptions 1, 3 and 5, for any $k\geq 0$, we have
$F_{k+1}\leq F_{k}-\frac{M_{k}-L}{4}R_{k}+\frac{\Delta_{k}}{M_{k}-L}.$
Remark 3.
The constants in Lemma 2 can be slightly improved if Assumption 1 is replaced by the convexity of $\psi$.
@@
$\leq f(\mathbf{x}_{k})+\langle\nabla f(\mathbf{x}_{k}),\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle$
$\quad+\frac{L}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+\psi(\mathbf{x}_{k+1})$
$=f(\mathbf{x}_{k})+\langle\mathbf{m}_{k},\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle+\psi(\mathbf{x}_{k+1})$
$\quad+\frac{M_{k}}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}-\frac{M_{k}-L}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}$
$\quad+\langle\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k},\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle\,.$
Then we apply Assumption 1 and notice that $\Omega_{k}(\mathbf{x}_{k})=\psi(\mathbf{x}_{k})+\langle\mathbf{m}_{k},\mathbf{x}_{k}\rangle$:
$F(\mathbf{x}_{k+1})$ $\leq F(\mathbf{x}_{k})-\frac{M_{k}-L}{2}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}$
$\quad+\langle\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k},\mathbf{x}_{k+1}-\mathbf{x}_{k}\rangle\,.$
@@
$F(\mathbf{x}_{k+1})$ $\leq F(\mathbf{x}_{k})-\frac{M_{k}-L}{4}\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}$
$\quad+\frac{\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}}{M_{k}-L}\,.$
@@
Lemma 4.
Under Assumptions 3 and 5, for any $k\geq 0$,
$(M_{k}^{2}+L^{2})R_{k}\geq\frac{1}{3}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]-\Delta_{k}.$
@@
We can split $\nabla F(\mathbf{x}_{k+1})$ in the following way:
@@
$=\mathbf{m}_{k}+\nabla\psi(\mathbf{x}_{k+1})+(\nabla f(\mathbf{x}_{k})-\mathbf{m}_{k})$
$\quad+(\nabla f(\mathbf{x}_{k+1})-\nabla f(\mathbf{x}_{k}))\,.$
Therefore,
@@
Now apply the stationarity condition (5) on the first term, and use Assumption 3 on the third term, we get:
$\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\leq 3(M_{k}^{2}+L^{2})\left\lVert\mathbf{x}_{k+1}-\mathbf{x}_{k}\right\rVert^{2}+3\left\lVert\mathbf{m}_{k}-\nabla f(\mathbf{x}_{k})\right\rVert^{2}\,.$
Rearranging and taking expectations, we get the claim. ∎
@@
Lemma 5.
Let Assumptions 1, 3, 4 and 5 hold, and let $a\coloneqq\frac{3}{8L}$, $\gamma_{k}\coloneqq\frac{3L}{M_{k}-L}$ and $M_{k}>4L$ for any $k\geq 0$. Then, for any $k\geq 0$,
$\Phi_{k+1}\leq\Phi_{k}-\frac{1}{48M_{k}}{\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{k+1})\right\rVert^{2}\right]+\frac{27L\sigma^{2}}{4M_{k}^{2}}.$ (7)
We have the following simple corollary for constant stepsize coefficients:
Corollary 6.
Let algorithm 1 be run for $K\geq 1$ iterations for solving problem (1) under Assumptions 3, 4 and 5, with constant coefficients $M_{k}=M=4L+\frac{3^{\nicefrac{{3}}{{2}}}}{2}\sqrt{\frac{KL\sigma^{2}}{\Phi_{0}}}$ and $\gamma_{k}=\frac{3L}{M-L}$ for any $10\leq k\leq K-$ 1, where $\Phi_{0}\coloneqq F(\mathbf{x}_{0})-F^{*}+\frac{3}{8L}{\mathbb{E}}\left[\left\lVert\mathbf{m}_{0}-\nabla f(\mathbf{x}_{0})\right\rVert^{2}\right]$. Then,
${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{t})\right\rVert^{2}\right]\leq 48(3^{\nicefrac{{3}}{{2}}})\sqrt{\frac{L\Phi_{0}\sigma^{2}}{K}}+\frac{192L\Phi_{0}}{K},$ (8)
where $t$ is chosen uniformly at random from $\{1,\ldots,K\}$.
It is also possible to use time-varying coefficients $M_{k}$ which do not require fixing the number of iterations in advance. However, in terms of the convergence rate, it incurs an extra logarithmic factor.
Corollary 7.
Consider algorithm 1 for solving problem (1) under Assumptions 3, 4 and 5 with coefficients $M_{k}=\max\bigl{\{}\sqrt{\frac{(k+1)L\sigma^{2}}{\Phi_{0}}},4L\bigr{\}}$, $\gamma_{k}=\frac{3L}{M_{k}-L}$ for any $k\geq 0$, where $\Phi_{0}\coloneqq F(\mathbf{x}_{0})-F^{*}+\frac{3}{8L}{\mathbb{E}}\left[\left\lVert\mathbf{m}_{0}-\nabla f(\mathbf{x}_{0})\right\rVert^{2}\right]$. Then, for any $k\geq 1$, we have
${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{t(k)})\right\rVert^{2}\right]\leq 372\ln(ek)\Bigl{(}\sqrt{\frac{L\Phi_{0}\sigma^{2}}{k}}+\frac{L\Phi_{0}}{k}\Bigr{)},$ (9)
where $t(k)$ is chosen randomly from $\{1,\ldots,k\}$ with probabilities $\Pr(t(k)=i)\propto\frac{1}{M_{i-1}}$, $i=1,\ldots,k$, and $e\coloneqq\exp(1)$.
Let us point out that using a random iterate as the output of the algorithm is standard in the literature (see, e.g., Rakhlin et al. (2012)) and can be efficiently implemented without fixing the number of iterations in advance. We discuss this more carefully in Appendix E.
@@ 5.2 Initialization and Convergence Guarantees
As we can see from corollaries 6 and 7, the convergence rate of algorithm 1 depends on $\Phi_{0}\coloneqq\mathcal{O}(F_{0}+\frac{1}{L}\Delta_{0})$, where $\Delta_{0}\coloneqq{\mathbb{E}}\left[\left\lVert\mathbf{m}_{0}-\nabla f(\mathbf{x}_{0})\right\rVert^{2}\right]={\mathbb{E}}\left[\left\lVert(1-\gamma_{-1})\mathbf{m}_{-1}+\gamma_{-1}\mathbf{g}_{0}-\nabla f(\mathbf{x}_{0})\right\rVert^{2}\right]$ depends on $\mathbf{m}_{-1}$. There are subtle differences in how $\mathbf{m}_{-1}$ can be initialized between the non-composite and composite cases.
@@
When $\psi\equiv 0$, i.e. $F\equiv f$, we set the initial momentum $\mathbf{m}_{-1}\coloneqq-\frac{\gamma_{-1}}{1-\gamma_{-1}}\mathbf{g}_{0}$, then $\Delta_{0}={\mathbb{E}}\left[\left\lVert\mathbf{m}_{0}-\nabla f(\mathbf{x}_{0})\right\rVert^{2}\right]=\left\lVert\nabla f(\mathbf{x}_{0})\right\rVert^{2}\leq 2LF_{0}$, where the last inequality follows from Assumption 3 and the fact that $F\equiv f$. Therefore, in the non-composite case, we can initialize the parameters such that $\Phi_{0}=\mathcal{O}(F_{0})$.
@@
The composite case is slightly trickier. We set $\mathbf{m}_{-1}\coloneqq\mathbf{g}_{0}$ and get that $\mathbf{m}_{0}=\mathbf{g}_{0}$. Hence $\Delta_{0}={\mathbb{E}}\left[\left\lVert\mathbf{g}_{0}-\nabla f(\mathbf{x}_{0})\right\rVert^{2}\right]\leq\sigma^{2}$. Therefore, $\Phi_{0}=\mathcal{O}(F_{0}+\frac{\sigma^{2}}{L})$. When $\sigma^{2}=\mathcal{O}(LF_{0})$, we get the same $\Phi_{0}=\mathcal{O}(F_{0})$ as in the non-composite case.
@@
In the case that $\sigma^{2}$ is much larger than $F_{0}$, if we have access to a constant size (not depending on the target error) mini-batch initially, then we can set $\mathbf{g}_{0}=\frac{1}{b_{0}}\sum_{i=1}^{b_{0}}\nabla f(\mathbf{x}_{0},\xi_{i})$ where $b_{0}\coloneqq\lceil\frac{\sigma^{2}}{LF_{0}}\rceil$, i.e. $\mathbf{g}_{0}$ is a mini-batch stochastic gradient of size $b_{0}$. Then we have $\Phi_{0}=\mathcal{O}(F_{0}+\frac{\sigma^{2}}{b_{0}L})=\mathcal{O}(F_{0})$, which is the same as in the non-composite case.
We have thus proved the following convergence guarantee for Algorithm 1.
Theorem 1.
Consider algorithm 1, as applied to solving problem (1) under Assumptions 3, 4 and 5, run for $K=\mathcal{O}\bigl{(}\frac{L\Phi_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{L\Phi_{0}}{\varepsilon}\bigr{)}$ iterations with constant coefficients $M_{k}=M=4L+\frac{3^{\nicefrac{{3}}{{2}}}}{2}\sqrt{\frac{KL\sigma^{2}}{\Phi_{0}}}$ and $\gamma_{k}=\frac{3L}{M-L}$ for any $10\leq k\leq K-$ 1, where $\Phi_{0}\coloneqq F_{0}+\frac{3}{8L}{\mathbb{E}}\left[\left\lVert\mathbf{m}_{0}-\nabla f(\mathbf{x}_{0})\right\rVert^{2}\right]$, $F_{0}\coloneqq F(\mathbf{x}_{0})-F^{*}$ and $\varepsilon>0$ is a given target error. Then, for the point $\mathbf{x}_{t}$ chosen uniformly at random from $\{\mathbf{x}_{1},\ldots,\mathbf{x}_{K}\}$ it holds that ${\mathbb{E}}\left[\left\lVert\nabla F(\mathbf{x}_{t})\right\rVert^{2}\right]\leq\varepsilon$.
If $\psi\equiv 0$, we can initialize $\mathbf{m}_{-1}$ in such a way that $K=\mathcal{O}\bigl{(}\frac{LF_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{LF_{0}}{\varepsilon}\bigr{)}$. Otherwise, we can initialize $\mathbf{m}_{-1}$ in such a way that $K=\mathcal{O}\bigl{(}\frac{LF_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{\sigma^{4}}{\varepsilon^{2}}+\frac{LF_{0}}{\varepsilon}+\frac{\sigma^{2}}{\varepsilon}\bigr{)}$. Further, when the initial mini-batch of size $\lceil\frac{\sigma^{2}}{LF_{0}}\rceil$ is allowed, we can initialize $\mathbf{m}_{-1}$ in such a way that $K=\mathcal{O}\bigl{(}\frac{LF_{0}\sigma^{2}}{\varepsilon^{2}}+\frac{LF_{0}}{\varepsilon}\bigr{)}$.
When an initial mini-batch is not allowed, the convergence rate has an extra $\mathcal{O}(\frac{\sigma^{4}}{\varepsilon^{2}}+\frac{\sigma^{2}}{\varepsilon})$ term that is not present in the non-composite case. One natural question is whether this is an artifact of our analysis or inherent to the problem and the algorithm. Note that the first step of the Algorithm 1 coincides with the vanilla stochastic proximal gradient method. Consider the lower bound construction in Proposition 1: when starting at $\mathbf{x}_{0}=0$, which is a stationary point such that $F_{0}=0$, one such step would incur an $\mathcal{O}(\sigma^{2})$ error. Therefore, in the non-composite case without the initial mini-batch, the convergence rate must have some term that only depends on $\sigma^{2}$. In other words, the extra terms seem unavoidable.